which system has no solution? (a) lines 1 and 2 (b) lines 1 and 3 (c) lines 1 and 4 (d) lines 2 and 3.
which system has infinite solution? (a) lines 1 and 2 (b) lines 1 and 3 (c) lines 1 and 4 (d) lines 2 and 4.
which system has exactly one solution? (a) lines 1 and 2 (b) lines 1 and 3 (c) line 2 and 4 (d) all of these
here are the equations
x+y=4
2x-3y=-9
y=2/3x
y=-x+4

Answer by stanbon(75887)

which system has no solution?
(a) lines 1 and 2
(b) lines 1 and 3
(c) lines 1 and 4
(d) lines 2 and 3.
Ans: d
--------------------------------------
which system has infinite solution?
(a) lines 1 and 2
(b) lines 1 and 3
(c) lines 1 and 4
(d) lines 2 and 4.
Ans: c
--------------------------------------
which system has exactly one solution?
(a) lines 1 and 2
(b) lines 1 and 3
(c) line 2 and 4
(d) all of these
Ans: d
--------------------------------------

here are the equations
x+y=4
slope = -1;intercept = 4
--------------------
2x-3y=-9
slope = 2/3 ; intercept = 3
--------------------
y=2/3x
slope = 2/3 ; intercept = 0
-----------------------
y=-x+4
slope = -1 ; intercept = 4
=================================
Cheers,
Stan H.